OpenStudy (anonymous):

The organic compound di-n-butyl phthalate, C16H22O4(l), is sometimes used as a low-density (1.046 g·mL–1) manometer fluid. Compute the pressure (in torr) of a gas that supports a 535-mm column of di-n-butyl phthalate. The density of mercury is 13.53 g·mL–1. The organic compound di-n-butyl phthalate, C16H22O4(l), is sometimes used as a low-density (1.046 g·mL–1) manometer fluid. Compute the pressure (in torr) of a gas that supports a 535-mm column of di-n-butyl phthalate. The density of mercury is 13.53 g·mL–1. @Chemistry

OpenStudy (aaronq):

The height ratio is the inverse of the density ratio. (d of 1/d of Hg) = (height Hg/height 1) (1.046/13.53) = (height Hg/555) (1.046/13.53) x 555 mm =

OpenStudy (aaronq):

1 mm Hg = 1 torr

OpenStudy (aaronq):

sorry you said 535 not 555